# Auto-Calibration and Biconvex Compressive Sensing with Applications to Parallel MRI

### v1

@@ 2.1 Related work in auto-calibration and compressive sensing
where $\mathbf{h}\in\mathbb{C}^{k}$, $\mathbf{y}\in\mathbb{C}^{L}$, $\mathbf{A}\in\mathbb{C}^{L\times N},L\leq N$ and $\|\mathbf{x}\|_{0}=n$. In this model, the unknown sensing parameters come after compression (e.g. matrix A is Gaussian or partial Fourier). Using the idea of lifting, Sparselift [36] first transforms the bi-convex constraint (6) to a linear constraint with respect to $\textbf{X}=\textbf{h}\textbf{x}^{T}$. Exploiting the sparsity in $\mathbf{x}$ and consequently within X, the recovery solves a convex relaxation by minimizing the $\|\mathbf{X}\|_{1}$ subject to the linear constraints. The authors show the recovery of $\mathbf{X}_{0}=\mathbf{h}_{0}\mathbf{x}_{0}^{T}$ is guaranteed with high probability when the sensing matrix $\mathbf{A}$ is either a Gaussian random matrix or with its rows chosen uniformly at random with replacement from the DFT matrix, when the number of measurements $L$ is of the order of $\mathcal{O}(kn\log^{2}(L))$. In a related work, [20] extends the results of SparseLift to the case where one observes $\mathbf{y}=\sum\limits_{i\in[r]}\text{diag}(\mathbf{B}\mathbf{h}_{i})\mathbf{A}\mathbf{x}_{i}+\boldsymbol{\omega}$. This extension improves recovery guarantees by solving a $l_{1,2}$ minimization by promoting block sparsity in the lifted signals. One important application of the auto-calibration problem in imaging includes reconstruction using randomly coded masks [6][51].
@@
On the other hand, parallel imaging (i.e., $C>1$ in (2)) consistently yields superior reconstruction results when compared to the single-coil ($C=1$) scenario [34][45][28]. We aim to establish a framework for comprehensively analyzing the role of multiple coils and the underlying reasons for its improved recovery.
From a practical viewpoint a relevant scenario focuses on self-calibration from multiple snapshots. For example, blind image restoration from multiple filters [26] and self-calibration model for sensors [7][35][22][9]. The goal here is to recover the unknown gains/phases $\mathbf{D}=\text{diag}(\mathbf{d})$ and a signal matrix $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{p}]$ from the measurement matrix $\mathbf{Y}=[\mathbf{y}_{1},\cdots,\mathbf{y}_{p}]$ and $\mathbf{Y}=\mathbf{D}\mathbf{A}\mathbf{X}$. In general, the method does not require sparse priors on the signal or the calibrations. Among this line of research, an interesting and somewhat related method involves using the idea of ‘linearization’ to transform the biconvex constraint into a linear one. One crucial assumption is that $\mathbf{s}$ does not have zero entries and define $\mathbf{d}=\mathbf{1/s}$, observations of the form $\mathbf{y}_{l}=\text{diag}(\mathbf{s})\mathbf{A}\mathbf{x_{l}}$ can be written as
$\mathbf{y}_{l}\mathbf{d}=\mathbf{A}\mathbf{x_{l}},$ (7)
which is a linear equation in $[\mathbf{s},\mathbf{x}_{l}]$. Compared to ‘Lifting’, the underlying problem dimension is unchanged. The same idea of linearization can also be found in [22][9][7][54]. In [35], the authors solve the problem using least-square minimization and give a theoretical recovery guarantee when $\mathcal{A}$ is generic.
For a general discussion on the issue of injectivity and the principle of identifiability for bi-linear compressive sensing, see e.g. [29][12]. It is also worth noting that [22][11] have developed self- calibration algorithms by using sparsity, convex optimization, and/or ideas from compressive sensing.
2.2 Related work in MRI
In this section, we show the connection of our problem to magnetic resonance imaging (MRI) and discuss auto-calibration methods within the specific application. MRI is a widely used imaging modality that enables high-contrast imaging. In a parallel MRI setting, multiple receiver coils are used to collect data from the same underlying image/3D volume. Data collected by the sensors are samples of the spatial Fourier transform of the underlying image multiplied by the complex receiver coil profile. We refer to the book [39] for an introduction to the physics of MRI. The need for precise calibrations of the coil profiles arises due to their non-uniformity, including:
  • Nonuniform sensitivity: coils are most sensitive near the body surface, diminishing with distance from the imaged object.
  • Phase modulation: coils induce point-wise phase shifts in the sensed signal.
The coil profile can be regarded as a complex vector $\mathbf{s}\in\mathbb{C}^{N}$ of the same length as the object to be imaged $\mathbf{x}\in\mathbb{R}^{N}$ and the forward model of pMRI can be written in the form of equation (1). The Fourier measurement (k space) is subsampled to achieve faster imaging.
MR images often give rise to sparse representations, either in the spatial domain or, more commonly, in some transform domain [37]. Since Lustig et al. [37] first demonstrated the compressive sensing approach for MRI reconstruction, it has become one of the most important tools used for modern MR imaging research. The redundancy in MRI can come in four ways: a) transform sparsity: natural images are sparse in transform domains, b) k-space (Fourier space) redundancy: exploit structured low-rankness of some formed measurements matrix which rely on the Fourier convolutional relationship, c) coil redundancy: sensitivity encoding by using multiple coils and their pairwise relationships, d) temporal redundancy: dynamic MRI. Coil domain redundancy uses the pairwise relationship between coils. The two most famous methods include SENSE [45] and GRAPPA [23] and their references within. Compared to the spatial/coil domain sparsity, methods exploiting structured k-space redundancy are relatively new. These approaches are derived from the k-space convolution relationship. We refer to SAKE [50], LORAKS[25], ALOHA [28] for this line of research. In this paper, we exploit both transform sparsity and coil redundancy. A similar setup, though in the perfectly calibrated case, is used in sparse SENSE [13] where the authors exploit joint sparsity across channels by solving a $l_{1,2}$ minimization problem.
From compressive sensing theory we know that by exploiting sparsity, under certain conditions images with a sparse representation can be recovered from randomly undersampled k-space (Fourier space) [10][49]. In our model, we assume our signal is sparse in some transform domain $\boldsymbol{\Psi}$. The choice of such $\boldsymbol{\Psi}$ is related to the field of research in sparsifying transformations [16][44]. More recently, attention has turned to adapting these models to data [3][48]. In this paper, we only consider a known sparsifying transformation $\mathbf{\Psi}$ satisfying $\mathbf{\Psi}^{*}\mathbf{\Psi}=\textbf{I}_{N}$. In the MRI literature, it has been demonstrated that the discrete cosine transformation, the wavelet transform and the finite difference transform are the mainly utilized choices for sparsifying MR images [50][47][13].
In practice, the coil profiles are often assumed to lie in a low dimensional subspace. Hence, B in (2) is commonly chosen as a polynomial basis or a sinusoidal basis as seen in practice [45] [24]. As for common practices for calibrating the coil profiles, typically, these profiles are estimated through a separate pre-scan. However, this approach is susceptible to motion errors, environmental fluctuations, and requires longer acquisition times, among other challenges. For auto-calibration methods, a line of research relies on a fully sampled k-space region [40][41] to extract the sensitivity parameters. Data driven auto-calibration methods, such as GRAPPA [23] and SPIRiT [38] can be viewed as interpolation methods by estimating linear relationships within the fully sampled k-space (e.g., kernel calibration) and enforce that relationship to synthesize data values in place of unacquired data (e.g., data reconstruction). More recently, a data-driven method SAKE [50] structures the multi-coil dataset into a new data matrix that is designed to have low-rankness property. Then it solves a matrix completion problem to fill in the missing data. Joint estimation techniques attempt to iteratively estimate both the coil sensitivities and image contents while imposing some smoothness constraints on the sensitivity profiles, some examples include [56][53]. Of particular interest is ENLIVE [27] where the author shows a connection between the original non-convex formulation to a nuclear norm minimization problem by using the idea of ‘Lifting’.
ENLIVE solves the following relaxed joint optimization problem over $k$ sets of coil profiles $\mathbf{s}^{i}_{j}$ and images $\mathbf{x}^{i}$ from measurements $\mathbf{y}_{i}$ over $C$ coils,
@@ 2.2 Related work in MRI
The operator $\mathbf{W}$ enforces smoothness in the coil profiles. Here, k sets of images $\mathbf{x}^{i}$ are used instead of one single image $\mathbf{x}$ to account for model violations. In a post-processing step, an average image is calculated as the solution.
The author shows an equivalence relationship between the formulation to a linearly constraint nuclear norm minimization problem by using the idea of ‘Lifting’. By defining the matrix $\mathbf{y}=\begin{bmatrix}\mathbf{y}_{1}&\cdots&\mathbf{y}_{C}\end{bmatrix}$, $\mathbf{U}=\begin{bmatrix}\mathbf{x}^{1}&\cdots\mathbf{x}^{k}\end{bmatrix}$, $\mathbf{V}=\begin{bmatrix}\mathbf{W}\mathbf{s}^{1}_{1}&\cdots\mathbf{W}\mathbf{s}^{k}_{j}\end{bmatrix}$, the problem is lifted into
@@
which is further related to the nuclear norm minimization problem with respect to the tensor product $\mathbf{Z}=\mathbf{U}\mathbf{V}^{T}$,
@@
ENLIVE is related to our approach as it uses lifting and exploits the low-rankness property of the lifted signal, whereas our method exploits the block-sparsity of the lifted signal. However, since nuclear norm minimization is computationally expensive, (10) is not utilized for practical computations but rather to provide an analytical explanation; In practice, the non-convex formulation (8) is used to solve the auto-calibrated reconstruction problem instead.
Although these algorithms undoubtedly offer utility and applicability by giving successful numerical results, the majority of them do not provide sufficient theoretical guarantees of recovery.

### v2

@@ 2.1 Related work in auto-calibration and compressive sensing
The SparseLift framework [ling_strohmer_2015] solves the auto-calibration problem in (7) by combining ‘lifting’ and bi-convex compressive sensing. The authors in [ling_strohmer_2015] first transform the biconvex constraint (7) to a linear constraint with respect to $\textbf{X}=\textbf{h}\textbf{x}^{T}$. Exploiting the sparsity in $\mathbf{x}$ and consequently within X, the recovery solves a convex relaxation by minimizing the $\|\mathbf{X}\|_{1}$ subject to the linear constraints. The authors show the recovery of $\mathbf{X}_{0}=\mathbf{h}_{0}\mathbf{x}_{0}^{T}$ is guaranteed with high probability when the sensing matrix $\mathbf{A}$ is either a Gaussian random matrix or with its rows chosen uniformly at random with replacement from the discrete Fourier transformation (DFT) matrix, when the number of measurements $L$ is of the order of $\mathcal{O}(kn\log^{2}(L))$. Note that the case where the signal is only approximately sparse is not analyzed. In a related work, the author in [flinth_2017] extends the results of SparseLift to the case where one observes a summation of such measurements from multiple sparse signals, $\mathbf{y}=\sum\limits_{i\in[r]}\text{diag}(\mathbf{B}\mathbf{h}_{i})\mathbf{A}\mathbf{x}_{i}+\boldsymbol{\omega}$. This extension improves recovery guarantees by solving a $\ell_{1,2}$ minimization by promoting block sparsity in the lifted signals.
The idea of ‘lifting’ also applies to another class of self-calibration problems, notably the blind deconvolution problem [Ahmed_Recht_Romberg_2014, Levin_Weiss_Durand_Freeman_2011]. In the general form of blind deconvolution, one measures the convolution of a signal with an unknown filter. That is, $\textbf{y}=I_{\Omega}(\textbf{s}*\textbf{x})+\textbf{n}$ where both s and x are unknown. Notice that the forward model for our problem in (2) can be written in the form of a convolution $\mathbf{y}=\mathbf{I}_{\Omega}(\mathbf{F}\mathbf{Bh}*\mathbf{F}\boldsymbol{\Psi^{*}}\mathbf{z})+\textbf{n}$ using the convolution theorem. When the Fourier space is fully sampled (i.e. $\mathbf{I}_{\Omega}=\mathbf{I}$ ), the problem is equivalent to recovery from
$\mathbf{F}^{-1}\mathbf{y}=\mathbf{Bh}*\boldsymbol{\Psi^{*}}\mathbf{z}+\mathbf{w},$ (8)
where $\mathbf{w}$ is also Gaussian. The work by Ahmed et al. [Ahmed_Recht_Romberg_2014] provides a theoretically guaranteed solution to the fully sampled case in (8). Using lifting, the authors solve a nuclear norm minimization problem by exploiting low-rankness in the lifted signal. If the active coefficients of z are known and $\|\textbf{z}\|_{0}=n$, the main theorem of [Ahmed_Recht_Romberg_2014] guarantees a stable recovery when the number of measurements is, up to log factor, of the order $\mathcal{O}(n+k)$ when $\boldsymbol{\Psi}$ is a random Gaussian matrix and B is incoherent. For the more difficult case when the convolution is only subsampled, the authors in [Lee_Li_Junge_Bresler_2017] provide rigorous analysis when both s and x are transform sparse (i.e., there exists $\boldsymbol{\Phi}$ and $\boldsymbol{\Theta}$ such that $\boldsymbol{\Phi}\textbf{x}$ and $\boldsymbol{\Theta}\textbf{s}$ are sparse). It provides an iterative algorithm for robust recovery guarantees with near optimal sample complexity. However, the recovery can only be guaranteed when both the sparsifying transformations (i.e. $\boldsymbol{\Phi}$ and $\boldsymbol{\Theta}$) are random Gaussians. As seen in many theories, subgaussian random matrices provide optimal measurement matrices for compressive sensing [foucart_rauhut_2015]. However, note that for our application case, $\boldsymbol{\Phi}=\boldsymbol{\Psi}\textbf{F}^{-1}$ and $\boldsymbol{\Theta}=\textbf{B}^{*}\textbf{F}^{-1}$, both impose structures due to physical constraints.
Similar to the parallel imaging (i.e., $C>1$ in (2)), from a practical viewpoint, a relevant scenario focuses on self-calibration from multiple snapshots. For example, blind image restoration from multiple filters [Harikumar_Bresler_1999] and self-calibration model for sensors [Balzano_Nowak_2007, Ling_Strohmer_2018, Gribonval_Chardon_Daudet_2012, Bilen_Puy_Gribonval_Daudet_2014]. The goal here is to recover the unknown gains/phases $\mathbf{S}=\text{diag}(\mathbf{s})$ and a signal matrix $\mathbf{X}=[\mathbf{x}_{1},\cdots,\mathbf{x}_{p}]$ from the measurement matrix $\mathbf{Y}=[\mathbf{y}_{1},\cdots,\mathbf{y}_{p}]$ and
$\mathbf{Y}=\mathbf{S}\mathbf{A}\mathbf{X}.$ (9)
For this model, the sensing matrix $\mathbf{A}$ is fixed through the sensing process, just as in the pMRI case. In general, the method does not require sparse priors on the signal or the calibrations. Among this line of research, an interesting and somewhat related method involves using the idea of ‘linearization’ to convexify the biconvex measurements (9) into a linear one. One crucial assumption for the ‘linearization’ scheme is that $\mathbf{s}$ does not have zero entries. Define $\mathbf{s}=\mathbf{1/d}$, the equations $\mathbf{y}_{l}=\text{diag}(\mathbf{s})\mathbf{A}\mathbf{x_{l}}$ can be written as
$\mathbf{y}_{l}\mathbf{d}=\mathbf{A}\mathbf{x_{l}},$ (10)
which is a linear equation in the concatenated vector $\begin{bmatrix}\mathbf{d}\\ \mathbf{x}_{l}\end{bmatrix}$. Compared to ‘lifting’, an advantage of ‘linearization’ is that the underlying problem dimension remains unchanged. The same idea of linearization can also be found in [Gribonval_Chardon_Daudet_2012, Bilen_Puy_Gribonval_Daudet_2014, Balzano_Nowak_2007, Wang_Chi_2016]. In [Ling_Strohmer_2018], the authors use ‘linearization’ and solve a least-square minimization problem for (9). Furthermore, theoretical recovery guarantee is provided when $\mathbf{A}$ is Gaussian.
As observant readers already may have noticed, the ‘true dimension’ of the problem of recovering $(\textbf{h}_{1},\cdots,\textbf{h}_{c},\textbf{z})$ is $(kC+n)$. To the best of the authors’ knowledge, currently there are no convex optimization schemes which succeed with such few measurements. In this context, the work by Oymak et al. [Oymak_Jalali_Fazel_Eldar_Hassibi_2015] should be mentioned. The authors demonstrated nonconvex procedures (e.g., minimizing a linear combination of $\ell_{0}$, $\ell_{0,2}$ or rank of the lifted matrix) that succeed with high probability, when the number of measurements is (up to log-factors) slightly larger than the information-theoretical limits. However, the theories apply only to measurement ensembles with specific properties, such as matrices with sub-Gaussian rows, subsampled standard basis (e.g. in matrix completion), or quadratic measurements for phase retrieval. Also, these procedures are not efficiently computable. For other structured measurement ensembles, it is worth noting the works by Qu et al. and Ling et al. [Qu_Li_Zhu_2019, Li_Ling_Strohmer_Wei_2019]. These studies address certain blind deconvolution problems with efficient algorithms and theoretical guarantees when the number of measurements is (up to log-factors) slightly larger than the information-theoretical limits. However, to the best of our knowledge, none of these methods directly apply to our measurement ensemble or require additional assumptions that do not align with our application case.
2.2 Developments in compressive sensing methods in pMRI
In the literature on pMRI, combining compressive sensing and parallel imaging has gained significant importance due to the need for fast acquisition. Accelerated imaging with fewer samples using multiple receiver coils has been an active field of research since its beginning. In this section, we briefly discuss the development of compressive sensing approaches in pMRI that exploit transform sparsity in the time domain, with a focus on those most closely related to our approach.
2.2.1 Relevant problem formulations and the optimization problems
In parallel imaging, the data are acquired from multiple channels simultaneously, but each at a rate lower than the Nyquist rate such that the data acquisition time is reduced. If only $1/R$ of the Fourier space is sampled per channel, the acquisition time is reduced by a factor of $R$.
The measurement model for pMRI using $C$ coils/channels is the following,
@@ 2.2.1 Relevant problem formulations and the optimization problems
where the encoding matrix E consists of the product of partial Fourier encoding with channel-specific sensitivity encoding over the image. Specifically, $\textbf{E}=\begin{bmatrix}\textbf{F}_{\Omega}\textbf{D}_{1}\\ \vdots\\ \textbf{F}_{\Omega}\textbf{D}_{C}\end{bmatrix}$ and $\textbf{D}_{i}=\text{diag}(\textbf{s}_{i})$ where $\textbf{s}_{i}$ is the coil sensitivity for the $i$-th coil.
According to linear algebra and the generalized sampling theorem, the maximum reduction factor is equal to the number of channels if E is of full rank. Because these conditions are rarely guaranteed in practice, the possible reduction factor is much lower. In practice, the reduction factor is typically chosen between 2 and 6 [Deshmane_Gulani_Griswold_Seiberlich_2012, Glockner_Hu_Stanley_Angelos_King_2005] for parallel imaging when the number of coils is large. Thus, at the boundary case, one would expect a threshold for the maximum reduction factor and the minimum number of channels necessary to ensure the invertibility of such a linear problem.
In addition to reducing sampling time through sensitivity encoding, further reduction in the number of samples can be achieved by exploiting redundancy in the signals. In this work, we specifically discuss time-domain sparsity, as magnetic resonance (MR) images typically exhibit sparse representation in some appropriate transform domain. Suppose it is assumed that the signal is $n$-sparse after some known transformation. One approach is to solve the following $\ell_{0}$ norm constrained least squares problem to directly exploit the $n$-sparsity of the signal:
$\min\limits_{\textbf{x}}\|\textbf{F}_{\Omega}\textbf{x}-\textbf{y}\|_{2}^{2}\quad\text{subject to }\|\boldsymbol{\Psi}\textbf{x}\|_{0}=n.$ (12)
As a convex relaxation to $\ell_{0}$ minimization, the following $\ell_{1}$ minimization is often utilized for practical compressive sensing MRI applications. Since Lustig et al. [lustig_donoho_pauly_2007] first demonstrated the compressive sensing approach for MRI reconstruction, it has become one of the most important tools used for modern MR imaging research. The original $\ell_{1}$ norm minimization problem solves the following convex program,
$\min\limits_{\textbf{x}}\|\boldsymbol{\Psi}\textbf{x}\|_{1}\quad\text{subject to }\textbf{F}_{\Omega}\textbf{x}=\textbf{y},$ (13)
where the signal x is assumed to be transform sparse. In the MRI literature, it has been demonstrated that the discrete cosine transformation, the wavelet transform, and the finite difference transform are the most commonly utilized choices for sparsifying MR images [shin_larson_ohliger_elad_pauly_vigneron_lustig_2014, Ravishankar_Bresler_2015, Chun_Adcock_Talavage_2014]. The choice of such $\boldsymbol{\Psi}$ is related to the field of research in sparsifying transformations [Elad_Milanfar_Rubinstein_2007, Pratt_Kane_Andrews_1969]. More recently, attention has turned to learning the sparsifying transformations from data [Aharon_Elad_Bruckstein_2006, Ravishankar_Bresler_2012].
In terms of combining CS with parallel imaging, the first applications were demonstrated in [Liang_Liu_Wang_Ying_2009, Bo_Liu_Yi_Ming_Zou_Leslie_Ying_2008, Ji_Chen_Zhao_Tao_Lang_2008, Wu_Millane_Watts_Bones_2010]. They can be seen as a direct extension of (13) to the parallel case,
$\min\limits_{\textbf{x}}\|\boldsymbol{\Psi}\textbf{x}\|_{1}+\alpha\|\textbf{x}\|_{TV}\quad\text{subject to }\textbf{E}\textbf{x}=\textbf{y},$ (14)
where $\textbf{D}_{i}=\text{diag}(\textbf{s}_{i})$ is the diagonal matrix of known calibration parameters of the $i$-th coil.
Note that both the convex relaxations in (13) (14) and nonconvex problem in (12) assume that the coil sensitivities are known. In practice, coil sensitivities are typically determined either through a separate pre-scan [pruessmann_weiger_scheidegger_boesiger_1999, mathematics_and_physics_of_emerging_biomedical_imaging_1996, Liang_Lauterbur_2000] or by interpolating from a fully sampled region of k-space (Fourier space) [griswold_jakob_heidemann_nittka_jellus_wang_kiefer_haase_2002, McKenzie_Yeh_Ohliger_Price_Sodickson_2002, Morrison_Jacob_Do_2007]. However, both methods have inherent drawbacks. The pre-scan approach is susceptible to motion artifacts, environmental fluctuations, and requires additional acquisition time. The interpolation method, while faster, may introduce errors due to imperfect k-space sampling or interpolation techniques. Moreover, both approaches may necessitate the use of filtering and kernel selection to obtain reasonable estimates of the coil sensitivities, which can lead to variations in the resulting reconstructions.
Given these challenges, auto-calibrated reconstruction methods have gained attention. In particular, it considers recovering all $\textbf{s}_{i}$ and x from the measurement model in (11). However, the linear system is highly underdetermined and ill-posed in this case. Indeed, if $\mathbf{s}_{i}$ and $\mathbf{x}$ form a pair of solutions to (11), then for any complex vector $\mathbf{v}$ with nonzero entries, $\mathbf{s}_{i}\odot\mathbf{v}$ and $\mathbf{x}\odot 1/\mathbf{v}$ form a pair of solutions.
Prior knowledge about the object and the coil sensitivities must be incorporated to render reasonable solutions. As formulated in Joint Image Reconstruction and Sensitivity Estimation in SENSE (JSENSE) [Ying_Sheng_2007] and Regularized Nonlinear Inversion (NLINV) [Uecker_Hohage_Block_Frahm_2008], the smoothness in the coil sensitivities is exploited. In JSENSE, the coil sensitivities are assumed to be expanded by a low-order polynomial basis (i.e. B) and the following bi-convex optimization problem is solved via alternating minimization,
$\arg\min\limits_{\mathbf{x},\mathbf{h_{j}}}\Sigma_{j=1}^{C}\|\mathbf{y}-\mathcal{F}_{\Omega}\{\textbf{B}\mathbf{h}_{j}\odot\mathbf{x}\}\|_{2}^{2}.$ (15)
In the NLINV formulation, they solve the following bi-convex optimization problem iteratively using a Gauss Newton type method,
$\arg\min\limits_{\mathbf{x},\mathbf{s_{j}}}\Sigma_{j=1}^{C}\|\mathbf{y}-\mathcal{F}_{\Omega}\{\mathbf{s}_{j}\odot\mathbf{x}\}\|_{2}^{2}+\alpha(\Sigma_{j=1}^{C}\|\mathbf{W}\mathbf{s}_{j}\|_{2}^{2}+\|\mathbf{x}\|_{2}^{2}),$ (16)
where $\mathbf{W}\in\mathbb{C}^{N\times N}$ is an invertible preconditioning matrix that penalizes high frequencies, thereby enforcing smoothness in the coil sensitivities. Other methods to enforce smoothness in the coil include expanding the coil sensitivities using low-order polynomial or sinusoidal basis functions, a method first introduced in the original SENSE paper [pruessmann_weiger_scheidegger_boesiger_1999] and later adopted in the analytical phantom package [Guerquin-Kern_Lejeune_Pruessmann_Unser_2012].
An extension of the NLINV (16) framework is the ENLIVE [Holme_Rosenzweig_Ong_Wilke_Lustig_Uecker_2019], which solves the following relaxed joint nonconvex optimization problem over $k$ sets of measurements,
@@
Here, k sets of images $\mathbf{x}^{i}$ are used instead of one single image $\mathbf{x}$ as in (16) to account for model violations. In a post-processing step, an average image using the $k$ sets of solutions is calculated as the solution.
Of particular interest to us is that ENLIVE can be related to a nuclear norm minimization problem that exploits the low-rankness in a ‘lifted’ domain. Specifically, by defining the matrix $\mathbf{y}=\begin{bmatrix}\mathbf{y}_{1}&\cdots&\mathbf{y}_{C}\end{bmatrix}$, $\mathbf{U}=\begin{bmatrix}\mathbf{x}^{1}&\cdots\mathbf{x}^{k}\end{bmatrix}\in\mathbb{C}^{N\times k}$, $\mathbf{V}=\begin{bmatrix}\mathbf{W}\mathbf{s}^{1}_{1}&\cdots\mathbf{W}\mathbf{s}^{k}_{1}\\ \vdots&\vdots&\\ \mathbf{W}\mathbf{s}^{1}_{C}&\cdots\mathbf{W}\mathbf{s}^{k}_{C}\end{bmatrix}\in\mathbb{C}^{CN\times k}$, the bi-linear measurements can be transformed into linear measurements in terms of a rank-k matrix $\textbf{U}\textbf{V}^{T}$:
@@
with a linear operator $\mathcal{A}$ that maps $\textbf{U}\textbf{V}^{T}$ to $\mathcal{F}_{\Omega}\{\Sigma_{i=1}^{k}\mathbf{s}_{j}^{i}\odot\mathbf{x}^{i}\}$. As shown in [Holme_Rosenzweig_Ong_Wilke_Lustig_Uecker_2019], the above formulation can be further relaxed to a nuclear norm minimization problem in terms of a matrix $\mathbf{Z}$,
@@
The authors claim that this connection provides an explanation for why ENLIVE favors low-rank solutions. However, it is worth noting that due to the computational cost associated with the nuclear norm minimization, the convex formulation (19) is not used for practical computations. Instead, the non-convex formulation (17) is employed to solve the auto-calibrated reconstruction problem. Notably, neither of the formulations in (15) nor (16) explicitly exploits sparsity in the object being imaged.
2.2.2 The relationship between $L$, $C$ and $n$
The primary objective of parallel imaging (PI) techniques is to reduce imaging time, which is achieved by acquiring fewer data points per coil and taking multiple coil measurements at the same time. From compressive sensing theory [candes_romberg_tao_2006, recht_fazel_parrilo_2010], it is well known that the incoherence in the sensing matrix is crucial for the successful solution of the problems in (13) and (14). And it is well accepted in the literature on pMRI that increasing the number of channels $C$ improves the incoherence of the sensing matrix E.
However, to understand the role of $C$ and $L$ versus the sparsity level $n$, a thorough analysis of the sensitivity encoding, partial Fourier sampling, and the sparsifying transformation in the overall measurement ensemble is required. As with linear inversion for solving (11), one would expect a threshold for the maximum reduction factor and a minimum number of coils for the compressive sensing approaches (e.g. $\ell_{0}$, $\ell_{1}$ or mixed norm minimization) to be successful. The problem becomes even more challenging, when the sensitivity encoding is unknown as in the auto-calibration framework.
In this context, the works by Otazo et al. [otazo2009distributed, Otazo_Kim_Axel_Sodickson_2010] should be mentioned,where they numerically assess the sufficient conditions on the number of measurements per coil $L$ and the number of coils $C$ for solving the $\ell_{0}$ norm minimization problem in (12) or the $\ell_{1}$ norm minimization problem in (13) to recover an $n$-sparse signal. The authors conclude that in the known calibration setup, when the number of coils is large enough, the required number of measurements per coil is very close to the signal sparsity for both the noiseless and noisy cases. To illustrate the numerical findings, we reproduced the reconstruction results from [otazo2009distributed] by solving the $\ell_{0}$ minimization problem (12), using the orthogonal matching pursuit (OMP) algorithm. The results can be found in Figure 1.
(a) Noise-free case
(b) With independent Gaussian noise
Figure 1: Two-dimensional pMRI simulations were performed, assuming truly transform n-sparse signals with randomly distributed non-zero components (N = 529, n = 32). Coils with sensitivities calculated using the Biot-Savart law were simulated for various numbers of coil elements ($n_{c}=2,4,6,8,12,16,24,36$). In the case of a single coil, a constant sensitivity was assumed. The multi-coil signal was generated by multiplying the signal with the coil sensitivities, with optional addition of Gaussian noise. The signal is assumed to be sparse after 2-dimensional DCT transforms. The average of the relative $\ell_{2}$ error was computed over 100 different random undersampling realizations in Fourier-space and random realizations of transform sparse signals, using varying numbers of measurements and coil elements.
At first glance, it may seem surprising that when $C$ is sufficiently large, increasing the number of coils does not yield any additional benefit in reducing the number of measurements per coil or the reconstruction quality. The following work may give us theoretical insights into the observation. Consider a more general setup in solving the following distributed compressed sensing problem:
$\mathbf{y}_{i}=\mathbf{F}_{\Omega_{i}}(\textbf{s}_{i}\odot\textbf{x}),i=1\cdots C,|\Omega_{i}|=L,$ (20)
where $\textbf{s}_{i}$ are known and $\mathbf{x}$ is n-sparse.
[…]
Note that this condition does not apply to multi-coil acquisition in pMRI, since all coils share the same undersampling pattern, albeit with distinct coil sensitivity profiles. Hence, one might expect a reduction in incoherence and a deviation from the theoretical recovery bounds discussed in [Duarte_Sarvotham_Baron_Wakin_Baraniuk]. Additionally, the measurement ensemble in the pMRI setting must adhere to additional physical constraints, which makes the analysis within the compressive sensing framework more complex. By combining the numerical experiments with the aforementioned theory for the known calibration case, one could reasonably expect that a sufficient condition for robust reconstruction would adopt a similar form to that of the best-case scenario for the more complex auto-calibration problem in (1).
2.2.3 Further discussions
Essentially, auto-calibration methods depend on data redundancy. In the literature on pMRI, other sources of redundancy have been utilized for auto-calibration, each under different assumptions. Data driven auto-calibration methods, such as GRAPPA [griswold_jakob_heidemann_nittka_jellus_wang_kiefer_haase_2002] and SPIRiT [Lustig_Pauly_2010] can be viewed as interpolation methods by estimating linear relationships within the fully sampled k-space (e.g., kernel calibration) and enforcing that relationship to synthesize data values in place of unacquired data (e.g., data reconstruction). More recently, a data-driven method SAKE [shin_larson_ohliger_elad_pauly_vigneron_lustig_2014] structures the multi-coil dataset into a new data matrix that is designed to have low-rankness property. Then it solves a matrix completion problem to fill in the missing data. Joint estimation techniques attempt to iteratively estimate both the coil sensitivities and image contents while imposing some smoothness constraints on the sensitivity profiles, some examples include [Ying_Sheng_2007, Uecker_Hohage_Block_Frahm_2008]. It is worth noting that the effectiveness of these methods depends on whether the coil and signal adhere to the assumptions made. As a result, they may produce varying quality of reconstructions for different datasets.